# SolNet: Open-source deep learning models for photovoltaic power forecasting across the globe

### v1

@@ Abstract
Deep learning models have gained prominence in the field of solar forecasting. One drawback of these models is that they require a lot of high-quality data. This is often not feasible, especially in solar power forecasting due to poor measurement infrastructure in legacy systems and the rapid build-up of new systems. This paper proposes SolNet: a novel, general-purpose solar power forecaster, which addresses these challenges by using transfer learning from abundant synthetic data generated from PVGIS. Using actual production data from hundreds of sites in Australia, Belgium and the Netherlands, we show that SolNet improves forecasting performance over data-scarce settings as well as simple benchmark models. We observe this effect to be the strongest when only limited data is available, and show that seasonal patterns have a major impact on the results. The study also quantifies synthetic data requirements to achieve these improvements, as well as the effect of misspecifications.
@@
Energy forecasting , Transfer learning , Neural Networks , Synthetic data , Seasonality
@@ 1 Introduction
Renewable energy sources (RES) are seen as a critical component of the energy transition necessary for climate change mitigation (panwar2011role). However, the actual generation from these RES depends on local weather conditions and is known for its variability and intermittency. As such, expanding generation through RES introduces additional uncertainty in the power system operation, necessitating sophisticated generation forecasts on hour-ahead to day-ahead levels (notton2018intermittent). However, with rapidly growing solar and wind installed capacity, both the average and worst case forecast errors have also rapidly grown. Solar power forecasting is especially problematic, with the worst-case day-ahead generation forecast error (mean absolute error calculated on hourly forecasts) having grown from around 3 GW to over 13 GW over 16 European countries (kazmi2022good). Moreover, the proliferation of rooftop systems in the distribution grid further complicates grid operation by introducing over-voltage and reverse power flow concerns (steffel2012integrating). Furthermore, unlike large solar plants where information is often freely available, forecasting power generation from small scale distributed solar systems is complicated by the lack of supporting forecasts for relevant exogenous factors (e.g. ambient conditions, cloud cover, etc.).
Motivated by these concerns, solar power forecasting has seen significant progress as a field in recent years (hong2020energy). Inspired by the state-of-the-art performance of deep learning in other domains such as computer vision and natural language processing, several researchers have also utilized these methods to build solar forecasters (kumari2021deep). The results of naively applying deep learning in forecasting, however, are mixed. Makridakis et al. (2018) talked about the seemingly under-performing pure deep learning models, compared to hybrid or pure statistical models after analysing the results of the M4 forecasting competition (makridakis2018m4), a key competition in the field of time series forecasting.
The Great Energy Predictor (GEP-III) competition, focused on forecasting building energy demand, resulted in similar conclusions about gradient boosting based methods outperforming deep learning methods (miller2020ashrae). A major reason for deep learning underperforming simpler models in many forecasting tasks is its requirement for enormous amounts of data to generalize well. On the contrary, conventional time-series models (e.g. those from the SARIMAX family), and simpler machine learning based models (e.g. those built on tree-based methods) can generalize relatively well from limited amounts of data. Consequently, they do not suffer from the fact that sufficient data on solar operation or exogenous factors is often not readily available in practice.
@@ 1.1 State-of-the-art
In recent years, transfer learning has emerged as a key ingredient to improve the sample complexity of deep learning methods (i.e. data required to generalise well) (weiss2016survey) in several domains including computer vision (CV) (gupta2022deep), natural language processing (NLP) (ruder-etal-2019-transfer), and reinforcement learning tasks (zhu2023transfer). In this scheme, a source model is trained initially on a large corpus of data (derived from a known source domain), which is then fine-tuned using sparse or limited observational data (derived from the target domain of interest). When building deep learning models from scratch, the initial pre-training phase (utilizing data from a source domain) is missing, leading to poor generalisation due to data sparsity. In practice, the closer the source domain is to the target domain, the better the results of transfer learning.
@@
One such recent example of transfer learning applied to PV power forecasting used a 3-layer LSTM architecture to predict PV power output, where the entire network was first trained on the source domain (solar irradiance), with the last layer fine-tuned on the target data (PV output) (zhou2020transfer). Interestingly, even though the source and target domain are somewhat different, the authors report improved predictive accuracy. Even more recently, abubakr2022unleashing have used transfer learning to predict solar irradiance at solar farms with limited data by using the abundant data available at one site. The model architecture was a Recurrent Neural Network (RNN), using historical solar irradiance data with up to four exogenous weather variables. This results in (slightly) better performance when using transfer learning compared to training the model directly on the limited data. The relatively small improvement in performance could well be due to the fact that the prediction time horizon is only one hour in the future.
@@
These studies are emblematic of the issues hampering adoption of transfer learning in the energy field. On the one hand, transferring from data observed at other, similar sites is not fully generalisable, i.e. it introduces a new requirement during the model’s training process. zhang2018generative used a generative adversarial network (GAN) to create synthetic time series of solar generation, which could theoretically improve forecasting accuracy. However, studies like this one are not common in literature to the best of our knowledge. They also suffer from some of the same issues: the GAN (or other trained generative model) requires previously collected datasets, which is the actual issue to begin with.
@@
The studies published so far also suffer from several limitations in terms of their evaluation robustness. They typically utilise very limited testing or future look-ahead periods, raising question marks about the statistical significance of the reported results. The seasonal variations of solar production over the year limits the generalisability of these results even more. Furthermore, they typically report improvements only in asymptotic performance, i.e. do not consider initial performance and learning performance. A more complete evaluation of transfer learning would look at differing amounts of training data, up until the initial performance, with no target data to fine-tune on. We elaborate on this further in the methodology section.
@@
Finally, arguably the most important reason behind the wide-spread adoption of source models in computer vision and natural language processing has been their availability to researchers and practitioners alike. The few source models that have been trained by researchers for energy or solar applications are all invariably closed-source and inaccessible, having been trained from scratch themselves (HU201683; QURESHI2017742; zhou2020transfer; abubakr2022unleashing). This largely defeats the purpose of transfer learning.
@@ 1.2 Contributions
Consequently, in this paper, we propose a novel way to forecast residential PV using deep-learning: SolNet, a unique open-source deep learning model generator for solar power forecasting for any land-based site on the planet. The proposed models, made available for use through a Python package (GitHub link in the conclusion section), rely only on API access to PVGIS (HULD20121803), to generate PV production data for several variations of weather conditions and panel specifications (e.g. tilt, azimuth, etc.). This removes the need for data collection for end-users, and makes the transfer approach more realistic. Furthermore, the models can be fine-tuned with observational data where it exists.
We compare the performance of SolNet to benchmarks using data for several years from 40 test locations in the Netherlands, 299 in Australia and 2 (commercial sites) in Belgium, respectively. We report all three performance metrics for better understanding of the value of transfer: initial performance, learning improvement, and asymptotic performance. Figure 1 visualises this idea. The initial performance of our models gives us the error when we have no (target) data to train on. For the target model this means that we use randomly initialised weights. For the transfer model this means that the model has pre-initialized weights based on the training that happened in the source domain. The learning performance is the increased performance by adding more data to the training process and the asymptotic performance is the difference in performance between the target and transfer model that we expect to have over time when enough data is added.
Figure 1: Complete assessment of transfer performance
@@ 1.3 Organization
The remainder of this paper is structured as follows: Section 2 discusses the transfer learning methodology and the models utilised for forecasting. Section 3 elaborates on the data used in this transfer learning process, including the source (PVGIS) and target (the actual site locations) domains, and how they are prepared for the analyses. Section 4 provides detailed results from applying the proposed methodology, while Sections 5 and 6 conclude the paper, going over the most important findings as well as some future research directions.
@@ 2 Methodology
We first discuss the design of the forecasting exercise, as well as the benchmark used for comparison purposes of our transfer approach. Next, we discuss the architecture of our chosen neural network, LSTM (long short term memory), which is a type of recurrent neural network which can learn temporal associations. Next, we elaborate on the transfer learning methodology applied to the LSTM models. Finally, we talk about the metrics used for the evaluation.
@@ 2.1 The forecasting exercise and baseline approach
For this paper we focused on the case of forecasting the next 24 hours, from 00:00AM to 23:00PM with no gap between inference time and the forecast period. This is a simple reference case to showcase the performance of SolNet, however the functionality is available for users to change the point of inference as well as the forecast period and it is our intention to benchmark the performance using different windows in later research.
To compare the performance of our deep learning model, we use a Naive seasonal model (K=24). This model takes the corresponding hour of the previous day as the prediction for the next day. So, if the actual value for 14:00 at day t was 0.45 kWh, the value for 14:00 at t+1 will also be 0.45 kWh. This method provides us with a realistic baseline, given the diurnal pattern of solar power. We will refer to this method as the ’naive’ or ’naive seasonal’ method for the remainder of this paper.
@@ 2.2 LSTM as a forecast model
The deep learning architecture used in our experiments, both with and without transfer learning, is the Long Short-term Memory (LSTM) model (hochreiter1997long). An LSTM model consists of several identical units. These units have several gates that remember (or forget) information over time. This allows the model to retain a form of memory, making it possible to learn long term patterns in time series data. All these gates are made from the fundamental building blocks of neural networks: weights, biases and non-linearities in the form of activation functions (e.g. sigmoid or tanh functions). These weights and biases are the parameters that are learned during the training process. The below formulas provide an overview of the internals of the LSTM unit.
@@
The relative simplicity of the LSTM model, existing out of N layers with identical building blocks, taking (scaled) inputs and running them through several activation functions after applying weights and biases to them, makes it a good candidate for transfer learning. The fine-tuning options are (relatively) limited compared to some modern architectures, and the only prerequisite of using the model in both the target domain and the source domain, is that the inputs to the models are of equal shapes (i.e. without introducing additional complexity). In the univariate case, this prerequisite is easily fulfilled as the model only requires the user to have K hours of historical PV output data available, with K being the number of historical lags used to make the forecast. For the purpose of this research we limit this to K=24. From trial and error we noted that adding more lags did not add significantly to training performance, whilst increasing the training time by a significant margin. Due to this research using over 300 test sites, this additional training time was not practical to account for.
The final architecture of the LSTM used in our experiments was found after doing a grid search for the best hyper-parameters (layers, neurons per layer, dropout percentage), and includes 3 layers, each consisting of 400 LSTM units, with one more layer consisting of 24 linear units. These 24 units form the outputs, one for each hour of the day. Other hyper-parameters include the optimizer used during the back-propagation process, we use ADAM based on existing literature, the learning rate, which is $10^{-4}$ in the base case and $10^{-6}$ in the fine-tuning case (see later in this section), a dropout rate to avoid overfitting (50% in our case), and a batch size (32). These parameters (learning rate, dropout rate and batch size) were chosen based on trial and error, as well as common useage in literature.
@@ 2.3 Transfer learning
With the deep learning architecture in place, it is possible to explore the effect of transfer learning on solar forecasting in a principled manner. More concretely, the LSTM can be pre-trained on source data, after which it can be fine-tuned using target data to achieve potential improvements over the target model. Transfer learning, formally defined as improving the learning of a target predictive function on the learning task, $\mathcal{F}_{T}(\cdot)$ in the target domain, $\mathcal{D}_{T}$, using knowledge of $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$ (the source domain and the source task) (pan2010survey). While $\mathcal{D}_{T}\neq\mathcal{D}_{S}$ and $\mathcal{T}_{T}\neq\mathcal{T}_{S}$, we do assume that, the more similar the two are, the easier the transfer learning task will be. In our case, we keep the source domain as closely related to the target domain as possible by mimicking the location and the PV set-up of our target domain data, in our source domain, but we also provide results for what happens when this condition does not hold (see section 4.4). Concretely, in addition to directly using the pre-trained source model for the target task of solar PV prediction, there are a few different algorithmic ways in which transfer learning can be applied to the LSTM networks:
@@
    Direct application. The pre-trained source model can be directly applied in the target domain without any adjustments to the model parameters.
@@
    Finetuning. The source LSTM can be fine-tuned to the target domain. We keep the architecture intact, and run several additional training iterations using the (limited) target domain data. We can avoid model overfitting on the limited target data by lowering the learning rate, thereby making sure that the model only makes slight adjustments to the previously trained weights. Another popular method to fine-tune deep neural networks is by ”freezing” the parameters of certain units, or entire layers (see e.g. 9070676). In this case, the optimiser can no longer update the parameters in these layers during the optimisation process.
For our experiments, we combine both finetuning strategies: we freeze the initial LSTM layer, and lower the learning rate of the subsequent layers by a factor of 100, when retraining on the target domain data. This ensures that the fine-tuning process does not completely over-write the learning from the source domain. We also look at direct application of the source models. This basically represents the case where we only evaluate the initial performance and no learning performance.

### v2

@@ Abstract
Deep learning models have gained increasing prominence in recent years in the field of solar photovoltaic (PV) forecasting. One drawback of these models is that they require a lot of high-quality data to perform well. This is often infeasible in practice, due to poor measurement infrastructure in legacy systems and the rapid build-up of new solar systems across the world. This paper proposes SolNet: a novel, general-purpose, multivariate solar power forecaster, which addresses these challenges by using a two-step forecasting pipeline which incorporates transfer learning from abundant synthetic data generated from PVGIS, before fine-tuning on observational data. Using actual production data from hundreds of sites in the Netherlands, Australia and Belgium, we show that SolNet improves forecasting performance over data-scarce settings as well as baseline models. We find transfer learning benefits to be the strongest when only limited observational data is available. At the same time we provide several guidelines and considerations for transfer learning practitioners, as our results show that weather data, seasonal patterns, amount of synthetic data and possible mis-specification in source location, can have a major impact on the results. The SolNet models created in this way are applicable for any land-based solar photovoltaic system across the planet where simulated and observed data can be combined to obtain improved forecasting capabilities.
@@
PV forecasting , Transfer learning , Synthetic data , Neural Networks , Seasonality
@@
[…]
[1]organization=ESAT, KU Leuven,addressline=Kasteelpark Arenberg 10, city=Leuven, postcode=3001, country=Belgium
@@ 1 Introduction
Renewable energy sources (RES) are seen as a critical component of the energy transition necessary for climate change mitigation (Panwar et al., 2011). However, the actual generation from these RES depends on local weather conditions and is known for its variability and intermittency. As such, expanding generation through RES introduces additional uncertainty in the power system operation, necessitating sophisticated generation forecasts on hour-ahead to day-ahead levels (Notton et al., 2018).
However, with rapidly growing solar and wind installed capacity, both the average and worst case forecast errors have also rapidly grown. Solar power forecasting is especially problematic, with the worst-case day-ahead generation forecast error (mean absolute error calculated on hourly forecasts) having grown from around 3 GW to over 13 GW in 16 European countries in just the past few years (Kazmi & Tao, 2022). Moreover, the proliferation of rooftop systems in the distribution grid further complicates grid operation by introducing over-voltage and reverse power flow concerns (Steffel et al., 2012). Additionally, unlike for large solar plants where information is systematically logged, forecasting power generation from these small scale distributed solar systems is further complicated by the lack of relevant meta-data (e.g. system size, orientation etc.) and supporting forecasts for related exogenous factors (e.g. ambient conditions, cloud cover, etc.).
Motivated by these concerns, solar power forecasting has seen significant progress as a field in recent years (Hong et al., 2020). Inspired by the state-of-the-art performance of deep learning in other domains such as computer vision and natural language processing, several researchers have also utilized these methods to build solar forecasters (Kumari & Toshniwal, 2021). The results of naively applying deep learning in forecasting, however, are mixed. Even though no solar-specific benchmarks are available, Makridakis et al. discuss the apparent under-performance of pure deep learning models, compared to hybrid or pure statistical models after analysing the results of the M4 forecasting competition (Makridakis et al., 2018), a key competition in the field of time series forecasting. The Great Energy Predictor (GEP-III) competition, focused on forecasting building energy demand, has likewise resulted in similar conclusions about simpler methods such as gradient boosting often outperforming deep learning methods (Miller et al., 2020).
A major reason for deep learning underperforming simpler models in many forecasting tasks is its requirement for enormous amounts of data to generalize well. On the contrary, conventional time-series models (e.g. those from the SARIMAX family), and simpler machine learning based models (e.g. those built on tree-based methods) can generalize relatively well from limited amounts of data. Consequently, these models do not suffer from the fact that sufficient historical data on solar operation or exogenous factors is often not readily available in practice.
@@ 1.1 State-of-the-art
In recent years, transfer learning has emerged as a key component in workflows aimed at improving the sample complexity of deep learning methods (i.e. the amount of data required to generalise well) (Weiss et al., 2016). This has been well-documented in several domains including computer vision (CV) (Gupta et al., 2022), natural language processing (NLP) (Ruder et al., 2019), and reinforcement learning tasks (Zhu et al., 2023). In this workflow, rather than training a model directly on (sparse) observational data, a source model is first trained on a large corpus of data (derived from a known source domain). This source model is then fine-tuned using sparse or limited observational data (derived from the target domain of interest). Foregoing the initial pre-training phase (i.e. utilizing data from the source domain) leads to poor generalisation due to data sparsity in practice. The closer the source domain is to the target domain, the better the results of transfer learning.
@@
Despite the pressing needs of the sector and the potential of transfer learning, practical examples of applying transfer learning to solar forecasting are still few and far between (Peirelinck et al., 2022). For instance, Zhou et al. have recently applied transfer learning to PV power forecasting used a 3-layer LSTM architecture to predict PV power output, where the entire network was first trained on the source domain (solar irradiance), with the last layer fine-tuned on the target data (PV output) (Zhou et al., 2020a). Interestingly, even though the source and target domain are somewhat different, the authors report improved predictive accuracy. Even more recently, Abubakr et al. (2022) have used transfer learning to predict solar irradiance at solar farms with limited data by using the abundant data available at one site. The model architecture was a Recurrent Neural Network (RNN), using historical solar irradiance data with up to four exogenous weather variables. This results in (slightly) better performance when using transfer learning compared to training the model directly on the limited data. The relatively small improvement in performance could well be due to the fact that the prediction time horizon is only one hour in the future.
@@
These studies are emblematic of the issues hampering adoption of transfer learning in the energy field. Transferring from data observed at other, similar sites is not fully generalisable, i.e. it introduces a new requirement during the model’s training process. One way to avoid this issue is by using synthetic data to perform the initial source model creation. However, despite early work by Zhang et al. using a generative adversarial network (GAN) to create synthetic time series of solar generation Zhang et al. (2018), this has not been pursued further. Using data-driven generative models also suffer from some of the same issues: the GAN (or other trained generative model) requires previously collected datasets to generalize well, which is the actual issue to begin with.
@@
The studies published so far also suffer from several limitations in terms of their evaluation robustness and transfer applicability. They typically utilise very limited testing or future look-ahead periods, raising questions about the statistical significance of the reported results. The seasonal variations of solar production over the year limits the generalisability of these results even more. Furthermore, these studies typically report improvements only in asymptotic performance, i.e. do not consider the influence of increasing data on model accuracy, a critical concern in transfer learning studies. A more complete evaluation of transfer learning would look at differing amounts of training data, up until the initial performance, with no target data to fine-tune on. We elaborate on this further in the methodology section.
@@
Finally, arguably the most important reason behind the wide-spread adoption of source models in computer vision and natural language processing has been their wide-spread availability to researchers and practitioners alike. The few source models that have been trained by researchers for energy or solar applications are all invariably closed-source and inaccessible, having been trained from scratch themselves (Hu et al., 2016; Qureshi et al., 2017; Zhou et al., 2020a; Abubakr et al., 2022). This largely defeats the purpose of transfer learning.
@@ 1.2 Contributions
Consequently, in this paper, we make two major contributions to the existing literature. First, we propose a novel way to forecast residential PV using transfer learning. This results in SolNet, a unique open-source deep learning model generator for solar power forecasting for any land-based site on the planet. The proposed models, made available for use through a Python package (GitHub link in section 6), rely only on API access to PVGIS (Huld et al., 2012) and Open-Meteo (Zippenfenig, 2023). PVGIS is used to generate PV production data for several variations of weather conditions and panel specifications (e.g. tilt, azimuth, etc.). Open-Meteo provides historical weather data, dating back for over 80 years. Using this large corpus of data for learning removes the need for data collection for end-users, makes the transfer approach more realistic, and applicable to largely any land-based location on the planet. Once the source model has been trained, it can be fine-tuned with observational data where it exists.
Second, we do a thorough deep-dive of various transfer learning scenarios, providing guidance on when and how to use it successfully in solar power generation scenario’s. We look at three performance metrics for better understanding of the value of transfer: initial performance, learning improvement, and asymptotic performance. Figure 1 visualises this idea. Additionally, we look at (1) the importance of weather data, (2) the impact of the yearly seasonality on transfer learning out-performance, and (3) how mis-specification of the source domain can impact the performance of transfer learning.
Figure 1: Complete assessment of transfer performance. The initial performance measures performance of a model without any training in the domain of interest (the target domain). The learning performance indicates the performance gain we can expect over time as more data to train on becomes available. Finally, the asymptotic performance is the ’permanent’ performance differential between the models when sufficient data is available.
@@ 1.3 Organization
The remainder of this paper is structured as follows: Section 2 discusses the forecasting exercise, the models utilised for forecasting, and the transfer learning methodology. Section 3 elaborates on the data used in this transfer learning process, including the source (PVGIS and Open-Meteo) and target (the actual site locations and DWD ICON EU) domains, and how they are prepared for the analysis. Section 4 provides detailed results from applying the proposed methodology, while Sections 5 and 6 conclude the paper, going over the most important findings as well as important future research directions.
@@ 2 Methodology
We first discuss the design of the forecasting exercise, as well as the benchmark used for comparison purposes of our transfer approach. Next, we discuss the architecture of our chosen neural network, LSTM (long short term memory), a type of recurrent neural network which can learn temporal associations. Then we elaborate on the transfer learning methodology applied to the LSTM models. Finally, we talk about the metrics used for the evaluation.
@@ 2.1 The forecasting exercise and baseline approach
Our forecasting exercise presents two cases, a univariate and multivariate case. In the univariate case we only use 24 lagged values of the target (PV power), and cyclically encoded values of the hour of the day. While we are aware that these are additional covariates, we still consider this case as univariate, as no external information is required for the cyclical encoding, we only rely on the hour of the day for these covariates. In the multivariate case we include several weather parameters as future covariates, as well as the 24 lagged values of PV power. The reader is referred to section 3.3 for more information on the weather data.
For this paper we focused on the case of forecasting the next 24 hours, from 00:00 to 23:00 with no gap between inference time and the forecast period. This is a simple reference case to showcase the performance of SolNet. The main reason for choosing this option is that, when using weather forecast data (which we get from the Deutscher Wetterdienst’s Icosahedral Nonhydrostatic EU model (DWD ICON EU) via Open Climate Fix (Open Climate Fix, 2023)), we have to decide on a reference timing. Historical weather forecasts, when and if they are available, typically offer the option to choose from four different reference time slots: 00:00, 06:00, 12:00 and 18:00. We limited ourselves to the 00:00 data, as otherwise the dimensionality of the data would have become too large to work with.
To compare the performance of our deep learning model, we use a Naive seasonal model (K=24). This model takes the corresponding hour of the previous day as the prediction for the next day. So, if the actual power generated at 14:00 on day $d$ was 0.45 kW, the forecast value for 14:00 on day $d+1$ will also be 0.45 kW. This method provides us with a realistic baseline, given the diurnal pattern of solar power. We will refer to this method as the ’naive’ or ’naive seasonal’ method for the remainder of this paper.
@@ 2.2 LSTM as a forecast model
The deep learning architecture used in our experiments, both with and without transfer learning, is the Long Short-term Memory (LSTM) model (Hochreiter & Schmidhuber, 1997). An LSTM model consists of several identical units. These units have several gates that remember (or forget) information over time, allowing the model to retain a form of memory. This enables LSTMs to learn long term patterns in time series data. All the gates in an LSTM are made from the fundamental building blocks of neural networks: weights, biases and non-linearities in the form of activation functions (e.g. sigmoid or tanh functions). These weights and biases are the parameters that are learned during the training process. The below formulas provide an overview of the internals of the LSTM unit.
@@
The relative simplicity of the LSTM model, existing out of N layers with identical building blocks, taking (scaled) inputs and running them through several activation functions after applying weights and biases to them, makes it a good candidate for transfer learning. The fine-tuning options are (relatively) limited compared to some modern architectures, and the only prerequisite of using the model in both the target domain and the source domain, is that the inputs to the models are of equal shapes (i.e. without introducing additional complexity). In the univariate case, this prerequisite is easily fulfilled as the model only requires the user to have K hours of historical PV output data available, with K being the number of historical lags used to make the forecast. For the purpose of this research we limit this to $K=24$. From trial and error we noted that adding more lags did not add significantly to training performance, whilst increasing the training time by a significant margin. Due to this research using over a hundred test sites with many configurations, this additional training time was not practical to account for.
The final architecture of the LSTM used in our experiments was found after doing a grid search for the best hyper-parameters (layers, neurons per layer, dropout percentage), and includes 3 layers, each consisting of 400 LSTM units, with one more layer consisting of 24 linear units. These 24 units form the outputs, one for each hour of the day. Other hyper-parameters include the optimizer used during the back-propagation process (ADAM), the learning rate ($10^{-4}$ in the base case and $10^{-6}$ in the fine-tuning case, see later in this section), a dropout rate to avoid overfitting (50% in our case), and batch size (32). These hyperparameters (optimizer, learning rate, dropout rate and batch size) were chosen based on trial and error and best practices from published literature.
@@ 2.3 Transfer learning
With the deep learning architecture in place, it is possible to explore the effect of transfer learning on solar forecasting in a principled manner. More concretely, the LSTM can be pre-trained on source data, after which it can be fine-tuned using target data to achieve potential improvements over the target model. Transfer learning, formally defined as improving the learning of a target predictive function on the learning task, $\mathcal{F}_{T}(\cdot)$ in the target domain, $\mathcal{D}_{T}$, using knowledge of $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$ (the source domain and the source task respectively) (Pan & Yang, 2010). While $\mathcal{D}_{T}\neq\mathcal{D}_{S}$ and $\mathcal{T}_{T}\neq\mathcal{T}_{S}$, we do assume that, the more similar the two are, the easier the transfer learning task will be. In our case, we keep the source domain as closely related to the target domain as possible by mimicking the location and the PV set-up of our target domain data, in our source domain, but we also provide results for what happens when this condition does not hold (see section 4.5). This is in addition to the inherent mis-alignment between simulated PVGIS data and observed data.
Concretely, in addition to directly using the pre-trained source model for the target task of solar PV prediction, there are a few different algorithmic ways in which transfer learning can be applied to the LSTM networks:
@@
    Direct application. The pre-trained source model can directly be applied in the target domain without any adjustments to the model parameters.
@@
    Finetuning. The source LSTM can be fine-tuned to the target domain. We keep the architecture intact, and run several additional training iterations using the (limited) target domain data. We can avoid model overfitting on the limited target data by lowering the learning rate, thereby making sure that the model only makes slight adjustments to the previously trained weights. Another popular method to fine-tune deep neural networks is by ”freezing” the parameters of certain units or entire layers (see e.g. Zhou et al. (2020b)). In this case, the optimiser can no longer update the parameters in these layers during the optimisation process, but is free to update the remaining parameters.
For our experiments, we combine both finetuning strategies: we freeze the initial LSTM layer, and lower the learning rate of the subsequent layers by a factor of 100, when retraining on the target domain data. This ensures that the fine-tuning process does not completely over-write the learning from the source domain. We also look at direct application of the source models. This basically represents the case where we only evaluate the initial performance (zero-shot learning).